# Learning Partial Differential Equations with Deep Parallel Neural Operator

### v2

@@ Abstract
In recent years, Solving partial differential equations has shifted the focus of traditional neural network studies from finite-dimensional Euclidean spaces to generalized functional spaces in research. A novel methodology is to learn an operator as a means of approximating the mapping between outputs. Currently, researchers have proposed a variety of operator architectures. Nevertheless, the majority of these architectures adopt an iterative update architecture, whereby a single operator is learned from the same function space. In practical physical science problems, the numerical solutions of partial differential equations are complex, and a serial single operator is unable to accurately approximate the intricate mapping between input and output. So, We propose a deep parallel operator model (DPNO) for efficiently and accurately solving partial differential equations. DPNO employs convolutional neural networks to extract local features and map data into distinct latent spaces. Designing a parallel block of double Fourier neural operators to solve the iterative error problem. DPNO approximates complex mappings between inputs and outputs by learning multiple operators in different potential spaces in parallel blocks. DPNO achieved the best performance on five of them, with an average improvement of 10.5%, and ranked second on one dataset.
@@
neural operator, scientific machine learning, deep learning-based PDE solver, neural network
journal: Neurocomputing
@@ 1 Introduction
The research in this paper focuses on developing a new operator architecture. The architecture uses Fast Fourier Transform (FFT) to learn partial differential equations. However, it solves the iterative error problem by using a clean and efficient parallel architecture. Our contributions are as follows:
@@
    We conduct six benchmark experiments to compare our model with benchmark models and currently recognized state-of-the-art models. Our model achieves the best results on five datasets with an average improvement of 10.5%, and ranks second on one dataset. In addition, we conduct an ablation study to compare the performance of parallel block with serial block. The results show an average performance improvement of 9.3% for parallel block.
@@ 2.2 Fourier transform and Fourier neural operator
$K(\phi)v_{t}(x)=\mathcal{F}^{-1}\left(\mathcal{R_{\phi}}\cdotp\left(\mathcal{F}v_{t}\right)(x)\right)$ (6)
2.3 Operator-learning Methods
@@
In the study of various partial differential equation problems, numerous operator design schemes have been proposed, yielding good results. Most operator-learning methods utilize convolutional neural networks and fully connected neural networks [24, 25, 26]. DeepONet, developed by Lu et al. [27], is based on the universal approximation theorem. The network comprises two parts, a branch network and a trunk network, which process the information from the sampled points of the input function and the target points, respectively. The outputs of these two parts are then combined to generate the final result, effectively approximating and predicting complex nonlinear operators. Geo-FNO works by transforming complex geometric tasks such as point clouds into a potentially unified mesh. Use this mesh as an input to FNO to efficiently process complex geometries [28]. F-FNO introduces new separable spectral layers and improved residual connections, combined with training strategies such as the Markov assumption, Gaussian noise, and cosine learning rate decay, significantly reducing errors on several benchmark PDEs [29]. MWT proposes a data-driven method for learning operator mappings by compressing operator kernels onto multi wavelet polynomial bases, utilizing the fine-grained representation of multi wavelet transforms and the properties of orthogonal polynomial bases [30]. SNO converts inputs and outputs into coefficients of basis functions, learning the mapping between inputs and outputs sequentially while alleviating systematic bias caused by aliasing errors [31]. KNO, based on Koopman theory, improves the performance of FNO in long-term predictions [32].
@@
In addition, there are other operator methods with deep architecture designs. Based on the U-Net concept and residual connections, multi-scale processing is performed for PDE tasks [33, 34]. For example , U-NO’s U-shaped structure improves model depth and memory efficiency. In benchmark tests for solving partial differential equations, such as Darcy’s and Navestock’s equations, U-NO significantly improves prediction accuracy [35]. U-FNO enhances the ability of FNO to handle multi-scale data by using residual connections [36]. LSM proposes a projection technique that sequentially projects high-dimensional data into multiple compact latent spaces, thereby eliminating redundant coordinate information. Based on triangular basis operators, LSM decomposes complex nonlinear operators into multiple basis operators through neural spectral blocks, thereby improving the ability to solve complex partial differential equations [37]. These novel neural operator learning methods have made significant advances and breakthroughs in different fields, demonstrating excellent performance.
@@ 3.1 Overall Architecture
Current mainstream neural operators usually adopt an iterative update strategy, which leads to the problem of spatial information loss due to iteration. To address this problem, we propose a parallel architecture based on hierarchical projection called Deep Parallel Neural Operator (DPNO). This approach maps the data to different latent spaces, approximates the complex mapping between these latent spaces, and thus fits the overall input-output mapping. We use several simple and efficient parallel operator blocks to learn the mapping relations within different latent spaces, by employing different combinations of modal truncations to deal with different frequency information.
@@
Overall architecture: Our model first maps data to different latent spaces through encoding, solves partial differential equations in multiple latent spaces, and finally returns to the original space through a decoder, as shown in Fig 1. Specifically, the DPNO architecture is divided into three main modules:
@@
where $\circ$ denotes the sequential combination between operations. In the DPNO architecture, $G_{\theta_{\text{toLatentSpace}}}$ denotes encoding data and mapping it to different latent spaces, and decoding $G_{\theta_{\text{toOriginalSpace}}}$ projects the data from latent space to original space. $G_{\theta_{\text{Ksolve}}}$ denotes a complex mapping that approximates inputs and outputs from latent space to different subspaces. The term “Loop” denotes a Loop. To improve model efficiency and reduce the number of parameters, $G_{\theta_{\text{Ksolve}}}$ shares the same set of parameters in the loop and learns the same set of parameters twice, i.e., the loop iterates twice.
@@ 3.2 Encoder
The coding network is a central part of DPNO. It maps data sequentially to multiple compact potential spatial representations. Consider that the Fourier transform is globally valid. Unlike simple downsampling methods, projection networks use classical convolutional combinations to extract local information from the data. Features are downsampled and compressed through pooling operations as a way to map to a lower dimensional space. This combination not only preserves key spatial information in the data, but also removes redundant noise. As a result, the effectiveness and expressiveness of the data representation can be enhanced while reducing the dimensionality. For the given input $x_{1}^{(L_{i})}$, the projection calculation is performed to obtain $x_{2}^{(L_{i})}:\mathcal{D}_{0}\rightarrow\mathcal{D}_{1}$. The combinatorial process is formally described as:
@@
3.3 Parallel operator block and Decoder
With a hierarchical projection network, we map the data to different potential spaces. We use a parallel dual Fourier neural operator to approximate the complex mapping between inputs and outputs. And a simple linear layer is used for residual connectivity to compensate for the loss of high-frequency information. The formal description is as follows:
@@ 4 Experiment and Results
In order to fully assess the performance of the model, we selected six partial differential equation benchmark datasets. The model relative errors were recorded. The selected benchmark datasets include typical partial differential equations in solid state physics and fluid physics and contain samples of different geometries. Details are given in Table 1.
@@ 4.1 Baselines and Settings
Table 1: Benchmark dataset.
@@
Table 2: Settings.
epoch batch_size learning_rate optimizer
500 20 0.001 Adam(1e-4)
To ensure that all models converge, we set the number of training rounds to 500. Training batch size of 20. Each parameter of the model will be updated using the Adam optimiser with a learning rate of 0.001 and a weight decay of 0.0001.
@@ 4.2.1 Navier-stokes Equation
The Navier Stokes equation is a fundamental equation that describes fluid motion. For viscous and incompressible fluids, we consider the famous Navier Stokes equation as a two-dimensional problem. The general form of the Navier Stokes momentum equation with periodic boundary conditions on a unit torus [0,1] is as follows:
$\frac{\partial u}{\partial t}+u\cdot\nabla u+\nabla p$ $=v\Delta u+f(x),\quad x\in T^{2},\ t\in(0,\infty)$ (14)
$\nabla u(x,t)$ $=0,\quad x\in T^{2},\ t\in[0,\infty)$
$u(x,0)$ $=u_{0}(x),\quad x\in T^{2}$
Where $\nu$ is the fixed viscosity, $u$ denotes the velocity, $p$ is the pressure, and $f$ is the external force function. The dataset is provided by the baseline model, and we used data with a grid resolution of $64\times 64$ for both training and testing. The results are shown in Table 3. Detailed visualisation results are in Appendix A.
Table 3: Comparison of Models for Navier-stokes Equation
@@ 4.2.2 Plastic Problem
The dataset is from the baseline model geo-FNO. Contains 980 training data and 80 test data. All data points are contained on a 101 × 31 structured grid with 20 time steps [28]. The results are shown in Table 4 and Fig 2.
(a) Ground True: Each column corresponds to a time point
(b) DPNO Predictions
(c) FNO Predictions
Figure 2: Plastic Problem
Table 4: Comparison of Models for Plasticity
@@ 4.2.3 Darcy Flow
Where $Q$ refers to the volume of fluid passing through the medium per unit time, $K$ is the permeability coefficient, $A$ is the flow cross-sectional area, and $\frac{dh}{dl}$ is the pressure gradient. We consider the steady state of the two-dimensional Darcy flow equation in a unit rectangle. It is a second-order, linear, elliptic partial differential equation with Dirichlet boundary conditions. We used the dataset provided in the baseline paper, with a grid size of 421 * 421. Other resolutions were extracted from the highest resolution, and our experiments used a fixed resolution of 85 * 85. The results are shown in Table 5 and Fig 3. To better demonstrate the generalization capability of the model, comparison experiments at different resolutions were added, as shown in Table 6.
Table 5: Comparison of Models for Darcy
@@
DPNO 0.0062(4.6%)
Table 6: Comparison of Models for Darcy at Different Resolutions
@@
(a) DPNO, FNO, and LSM predictions.
(b) Differences in absolute values between the predictions of DPNO, FNO, LSM, and the ground truth.
Figure 3: Darcy Flow.
@@ 4.2.4 Elastic Problem
where $I_{1}$ and $I_{2}$ are scalar invariants of the right Cauchy Green stretch tensor $C=2\epsilon+I$.$C_{1}$ is a tunable energy density parameter. In the experiments, we adopted 1000 training points and 200 testing points [28]. The results are shown in Table 7 and Fig 4.
Table 7: Comparison of Models for Elastic-G
@@
DPNO 0.0375(8%)
(a) DPNO, FNO, and LSM predictions.
(b) Differences in absolute values between the predictions of DPNO, FNO, LSM, and the ground truth.
Figure 4: Elastic Problem.
4.2.5 Airfoil problem with Euler’s equation
@@
where $\rho$ is the mass per unit volume, reflecting the concentration of the fluid; $\mathbf{u}$ is the distance travelled by a fluid particle per unit time, a vector describing the change in fluid momentum and flow characteristics; $p$ is the force acting per unit area; and $E$ is the total energy of the system. In this case, the far-field boundary conditions are $\rho_{1}=1$, $p_{1}=1.0$, $M_{1}=0.8$, and $\text{AoA}=0$, where $M_{1}$ is the Mach number and AoA is the angle of attack. Additionally, a no-penetration condition is imposed at the airfoil [28]. The experiment uses 1000 training grid points and 200 test grid points. The results are shown in Table 8 and Fig 5.
(a) DPNO, FNO, and LSM predictions.
(b) Differences in absolute values between the predictions of DPNO, FNO, LSM, and the ground truth.
Figure 5: Airfoil problem.
Table 8: Comparison of Models for Airfoil Problem

### v3

@@ Abstract
In recent years, Solving partial differential equations has shifted the focus of traditional neural network studies from finite-dimensional Euclidean spaces to generalized functional spaces in research. A novel methodology is to learn an operator as a means of approximating the mapping between outputs. Currently, researchers have proposed a variety of operator architectures. Nevertheless, the majority of these architectures adopt an iterative update architecture, whereby a single operator is learned from the same function space. In practical physical science problems, the numerical solutions of partial differential equations are complex, and a serial single operator is unable to accurately approximate the intricate mapping between input and output. So, We propose a deep parallel operator model (DPNO) for efficiently and accurately solving partial differential equations. DPNO employs convolutional neural networks to extract local features and map data into distinct latent spaces. Designing a parallel block of double Fourier neural operators to solve the iterative error problem. DPNO approximates complex mappings between inputs and outputs by learning multiple operators in different potential spaces in parallel blocks. DPNO achieved the best performance on five of them, with an average improvement of 11.6%, and ranked second on one dataset.
@@
Mechanical partial differential equations, neural operator, scientific machine learning, deep learning-based PDE solver
journal: Engineering Applications of Artificial Intelligence
@@ 1 Introduction
For several classical partial differential equations in fluid mechanics, The research in this paper focuses on developing a new operator architecture. The architecture uses Fast Fourier Transform (FFT) to learn partial differential equations. However, it solves the iterative error problem by using a clean and efficient parallel architecture. Our contributions are as follows:
@@
    We conduct six benchmark experiments to compare our model with benchmark models and currently recognized state-of-the-art models. Our model achieves the best results on five datasets with an average improvement of 11.6%, and ranks second on one dataset. In addition, we conduct an ablation study to compare the performance of parallel block with serial block. The results show an average performance improvement of 10.8% for parallel block.
@@ 2.2 Fourier transform and Fourier neural operator
$K(\phi)v_{t}(x)=\mathcal{F}^{-1}\left(\mathcal{R}_{\phi}\cdotp\left(\mathcal{F}v_{t}\right)(x)\right)\quad\forall x\in D$ (6)
2.3 Operator Learning Methods
@@
In the study of various partial differential equation problems, numerous operator design schemes have been proposed, yielding good results. Most operator-learning methods utilize convolutional neural networks and fully connected neural networks [24, 25, 26]. DeepONet, developed by Lu et al. [27], is based on the universal approximation theorem. The network comprises two parts, a branch network and a trunk network, which process the information from the sampled points of the input function and the target points, respectively. The outputs of these two parts are then combined to generate the final result, effectively approximating and predicting complex nonlinear operators. For the irregular grid problem, Li et al. [28] proposed the Geo-FNO model, Geo-FNO works by transforming complex geometric tasks such as point clouds into a potentially unified mesh. Use this mesh as an input to FNO to efficiently process complex geometries . Tran et al. [29] proposed a decomposed Fourier neural operator (F-FNO), F-FNO introduces new separable spectral layers and improved residual connections, combined with training strategies such as the Markov assumption, Gaussian noise, and cosine learning rate decay, significantly reducing errors on several benchmark PDEs. Gupta et al. [30] proposed the Multi-Wavelet Transform Neural Operator (MWT), MWT takes advantage of the fine-grained representation of the Multi-Wavelet Transform and the orthogonal polynomial bases by compressing the kernel of the operator onto a Multi-Wavelet polynomial basis. A strong approximation capability is demonstrated in several PDE tasks. Zhang et al. [31] proposed a spectral neural operator(SNO), SNO converts inputs and outputs into coefficients of basis functions, learning the mapping between inputs and outputs sequentially while alleviating systematic bias caused by aliasing errors . Xiong et al. [32] proposed the Koopman Neural Operator (KNO) based on the Koopman theory. The KNO replaces the feature mapping in the Fourier Neural Operator with the Koopman Operator, which improves the performance of the FNO in long-term prediction.
@@
In addition, there are other operator methods with deep architecture designs. Based on the U-Net concept and residual connections, multi-scale processing is performed for PDE tasks [33, 34]. For example, Zhang et al. [35] designed a U-shaped neural operator (U-NO), U-NO’s U-shaped structure improves model depth and memory efficiency. In benchmark tests for solving partial differential equations, such as Darcy’s and Navestock’s equations, U-NO significantly improves prediction accuracy. Wen et al. [36] frame-integrated FNO with U-Net to design an enhanced Fourier neural operator (U-FNO), U-FNO enhances the ability of FNO to handle multi-scale data by using residual connections. Wu et al. [37] proposed the latent spectrum model (LSM), LSM introduces a projection technique that sequentially projects high-dimensional data into multiple compact latent spaces, thereby eliminating redundant coordinate information. Based on triangular basis operators, the LSM decomposes complex nonlinear operators into multiple basis operators through neural spectral blocks, thereby enhancing the ability to solve complex partial differential equations.These novel neural operator learning methods have made significant advances and breakthroughs in various fields, demonstrating excellent performance.
@@ 3 Method
3.1 Problem Setup
For a PDE-governed task, consider a bounded open set $D\subset\mathbb{R}^{d}$ representing the coordinate domain. For this problem, inputs and outputs can be reformulated as functions defined over the coordinate domain, which belong to the Banach spaces $X=X(D;\mathbb{R}^{dx})$ and $Y=Y(D;\mathbb{R}^{dy})$, where $\mathbb{R}^{dx}$ and $\mathbb{R}^{dy}$ denote the ranges of the input and output functions, respectively.
At each coordinate $s\in D$, the function values $x(s)\in\mathbb{R}^{dx}$ and $y(s)\in\mathbb{R}^{dy}$ represent the input and output values at location $s$, akin to pixel values. Under this framework, the PDE-solving process is framed as approximating the optimal mapping operator $G:X\to Y$, typically using a deep neural network model $G_{\theta}$ to approximate this mapping, where $\theta\in\Theta$ represents the parameter set of the model. The model learns from observed samples $\{(x,y)\}$ to approximate $G$.
@@ 3.2 Overall Architecture
Current mainstream neural operators usually adopt an iterative update strategy, which leads to the problem of spatial information loss due to iteration. To address this problem, we propose a parallel architecture based on hierarchical projection called Deep Parallel Neural Operator. This approach maps the data to different latent spaces, approximates the complex mapping between these latent spaces, and thus fits the overall input-output mapping. We use multiple simple and efficient parallel Fourier layer modules to learn the mapping relationships within different latent spaces by employing different combinations of modal truncations to process different frequency information.
@@
Overall architecture: start from input $\alpha(x)$, Lift to a higher dimension channel space by a neural network P, Local feature extraction via $\theta$. The input data is mapped to the potential space by two projections, Learning the solution operator via a block of parallel Fourier layers and finally returns to the original space through a decoder, as shown in Fig 1. Specifically, the DPNO architecture is divided into three main modules:
@@
where $\circ$ denotes the sequential combination between operations. In the DPNO architecture, $G_{\theta_{\text{toLatentSpace}}}$ denotes encoding data and mapping it to different latent spaces, and decoding $G_{\theta_{\text{toOriginalSpace}}}$ projects the data from latent space to original space. $G_{\theta_{\text{Ksolve}}}$ denotes a complex mapping that approximates inputs and outputs from latent space to different subspaces. “Loop” denotes a Loop. To improve model efficiency and reduce the number of parameters, $G_{\theta_{\text{Ksolve}}}$ shares the same set of parameters in the loop and learns the same set of parameters twice, i.e., the loop iterates twice.
@@ 3.3 Encoder
The encoding network is a central part of DPNO. It maps data sequentially to multiple compact potential spatial representations. Consider that the Fourier transform is globally valid. Unlike simple downsampling methods, projection networks use classical convolutional combinations to extract local information from the data. Features are downsampled and compressed through pooling operations as a way to map to a lower dimensional space. This combination not only preserves key spatial information in the data, but also removes redundant noise. As a result, the effectiveness and expressiveness of the data representation can be enhanced while reducing the dimensionality. For the given input $x_{1}^{(L_{i})}$, the projection calculation is performed to obtain $x_{2}^{(L_{i})}:\mathcal{D}_{0}\rightarrow\mathcal{D}_{1}$. The combinatorial process is formally described as:
@@
3.4 Parallel Fourier layer block and Decoder
With a hierarchical projection network, we map the data to different potential spaces. We use a parallel double Fourier layer to approximate the complex mapping between input and output, and since modal truncation is performed after the Fourier transform, a simple convolutional layer is used via residual concatenation to compensate for the loss of high-frequency information. The formal description is as follows:
@@ 4 Experiment and Results
In order to fully assess the performance of the model, we selected six partial differential equation benchmark datasets. The model relative errors were recorded. In order to reflect the superiority and generalisation ability of the model, we have selected four partial differential equations in fluid mechanics in addition to two partial differential equations related to solid mechanics. Details are given in Table 1.
@@ 4.1 Baselines and Settings
Table 1: Summary of experiment benchmarks.
@@
Table 2: Experimental Setup.
epoch learning_rate optimizer
500 0.001 Adam(1e-4)
To ensure that all models converge, we set the number of training rounds to 500. Each parameter of the model will be updated using the Adam optimiser with a learning rate of 0.001 and a weight decay of 0.0001.
@@ 4.2.1 Navier-stokes Equation
The Navier Stokes equation is a fundamental equation that describes fluid motion. Consider the two-dimensional Navier-Stokes square for a viscous, incompressible fluid on the unit ring surface, which can be described in vorticity form as:
$\frac{\partial\omega(x,t)}{\partial t}+u(x,t)\cdot\nabla\omega(x,t)$ $=\nu\Delta\omega(x,t)+f(x),\quad x\in(0,1)^{2},\ t\in(0,T]$ (14)
$\nabla\cdot u(x,t)$ $=0,\quad x\in(0,1)^{2},\ t\in(0,T]$
$\omega(x,0)$ $=\omega_{0}(x),\quad x\in(0,1)^{2}$
Among them, $u\in C([0,T];H^{r}_{\text{per}}((0,1)^{2};\mathbb{R}^{2}))$ for any $r>0$ is the velocity field, $\omega$ is the vorticity, $\omega_{0}\in L^{2}_{\text{per}}((0,1)^{2};\mathbb{R})$ is the initial vorticity, $\nu>0$ is the viscosity coefficient, $f\in L^{2}_{\text{per}}((0,1)^{2};\mathbb{R})$ is the forcing function. The main purpose of the neural network is to learn operators that map vorticity before time 10 to vorticity at some later time $T>10$, $G:C([0,10];H^{r}_{\text{per}}((0,1)^{2};\mathbb{R}))\to C((10,T];H^{r}_{\text{per}}((0,1)^{2};\mathbb{R}))$, defined as $\omega:(0,1)^{2}$ from $[0,10]$ to $(10,T]$.
The initial condition $w_{0}(x)$ is derived from $w_{0}$, defined as
$w_{0}=N\left(0.73\cdot 2\left(\cos(x_{1}+x_{2})+4.9i\right),0.25\right),$
which satisfies periodic boundary conditions. The external forcing term is fixed as
$f(x)=0.1\left(\sin\left(2(x_{1}+x_{2})\right)+\cos\left(2(x_{1}+x_{2})\right)\right).$
The equation is solved using a stream function representation combined with a pseudospectral approach. Initially, a Poisson equation is resolved in the Fourier domain to determine the velocity field. Subsequently, the vorticity is differentiated, and nonlinear terms are computed within the physical domain, followed by a dealiasing process to eliminate high-frequency noise. Time integration employs the Crank-Nicolson scheme, treating the linear terms implicitly while keeping the nonlinear terms explicit. For data generation, downsampling is performed to obtain a $64\times 64$ grid resolution for training and testing. The results are shown in Table 3. Experimental results are plotted in A.
Table 3: Model performance comparison on the Navier-Stokes benchmark.
@@ 4.2.2 Plastic Problem
For this problem, we view the problem as a three-dimensional problem containing time. For neural networks, the main concern is to model the changing state of the material at a future moment through the initial state. The dataset is from the baseline model geo-FNO. Contains 980 training data and 80 test data. All data points are contained on a 101 × 31 structured grid with 20 time steps [28]. The results are shown in Table 4 and Fig 2.
Figure 2: Plastic Problem: The first row is the true value, rows 2-4 are the model error, and on the right side are the true value legend and the error legend, respectively. Each column represents a state at a point in time.
Table 4: Model performance comparison on the Plasticity problem benchmark.
@@ 4.2.3 Darcy Flow
Where $Q$ refers to the volume of fluid passing through the medium per unit time, $K$ is the permeability coefficient, $A$ is the flow cross-sectional area, and $\frac{dh}{dl}$ is the pressure gradient. We consider the steady state of the two-dimensional Darcy flow equation in a unit rectangle. It is a second-order, linear, elliptic partial differential equation with Dirichlet boundary conditions. The format is as follows:
$-\nabla\cdot\left(a(x)\nabla u(x)\right)$ $=f(x),\quad x\in(0,1)^{2},$ (17)
$u(x)$ $=0,\quad x\in\partial(0,1)^{2}.$
where $a\in L((0,1)^{2};\mathbb{R}^{+})$ is the diffusion coefficient and $f\in L^{2}((0,1)^{2};\mathbb{R})$ is the forcing function.
For the above equation, with respect to neural networks, our main focus is on how to obtain the solution operator through the diffusion coefficient. We used the dataset provided in the baseline paper, with a grid size of 421 * 421. Other resolutions were extracted from the highest resolution, and our experiments used a fixed resolution of 85 * 85. The results are shown in Table 5 and Fig 3. To better demonstrate the generalization capability of the model, comparison experiments at different resolutions were added, as shown in Table 6.
Table 5: Model performance comparison on the Darcy Flow benchmark under 85*85 resolutions.
@@
DPNO 0.0057
Table 6: Model performance comparison on the Darcy Flow benchmark under different resolutions.
@@
Figure 3: Darcy flow: The first row shows the true values and model predictions, the second row shows the model errors, and the legends are the true value legend and the error legend, respectively.
@@ 4.2.4 Elastic Problem
where $I_{1}$ and $I_{2}$ are scalar invariants of the right Cauchy Green stretch tensor $C=2\epsilon+I$.$C_{1}$ is a tunable energy density parameter. The dataset was generated by Li et al. [28] using a finite element solver with 1000 training data and 200 test data with approximately 100 quadratic quadrilateral elements. The input $\alpha$ is given in the form of a point cloud of approximately 1000 size. Since our neural operator network is only capable of receiving uniform rectangular grid data, uniform grid data is obtained by differencing the data. The neural network mainly fits a solution operator which is able to fit the strain of the material through the initial state of the material. In the experiments, we adopted 1000 training points and 200 testing points. The results are shown in Table 7 and Fig 4.
Table 7: Model performance comparison on the Elastic Problem benchmark.
@@
DPNO 0.0375
Figure 4: Elastic Problem: The first row shows the true values and model predictions, the second row shows the model errors, and the legends are the true value legend and the error legend, respectively.
4.2.5 Airfoil problem with Euler’s equation.
@@
where $\rho$ is the mass per unit volume, reflecting the concentration of the fluid; $\mathbf{u}$ is the distance travelled by a fluid particle per unit time, a vector describing the change in fluid momentum and flow characteristics; $p$ is the force acting per unit area; and $E$ is the total energy of the system. In this case, the far-field boundary conditions are $\rho_{1}=1$, $p_{1}=1.0$, $M_{1}=0.8$, and $\text{AoA}=0$, where $M_{1}$ is the Mach number and AoA is the angle of attack. Additionally, a no-penetration condition is imposed at the airfoil. The dataset was generated from 1000 training data and 200 test data using a second-order implicit finite volume solver by Li et al. A C-mesh with about 200 quadrilateral elements was used and adjusted for the mesh near the wing but not around the impact [28]. The neural network fits a mapping from the shape parameters of the airfoil to the flow field around the airfoil. The mesh point locations and Mach number on these mesh points are used as input and output data. The results are shown in Table 8 and Fig 5.
Figure 5: Airfoil problem: The first row shows the true values and model predictions, the second row shows the model errors, and the legends are the true value legend and the error legend, respectively.
Table 8: Model performance comparison on the Airfoil Problem benchmark.